# Deciphering interventional dynamical causality from non-intervention systems

### v1

@@ Methods
We introduce the IEE criterion, specifically designed for quantifying IntDC solely from the observational data in the delay-embedding space. The IEE criterion is proposed to measure the average information retention or reconstructability from the effect $\delta\bm{Y}_{t+1}$ to the causation $\delta\bm{X}_{t}$, formulated as
$\text{IEE}[x\rightarrow y]=\text{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1}),$ (6)
where CMI is the conditional mutual information  (?), i.e.
@@
The $\delta\bm{Y}_{t+1}$ is estimated by the distance from $\bm{Y}_{t+1}$ to its neighbored points $\bm{Y}_{t_{k}+1}$ in the embedding space, and $\delta\bm{X}_{t}$ is calculated by $\bm{X}_{t_{k}}$ with the same time label $t_{k}$ (Figs. 1(E) and (F)). Thus, Eq. (6) can be numerically approximated only from the observational data, especially for non-intervention systems. Table 1 shows the numerical algorithm for IEE, whose detailed derivation can be found in the Supplementary Text.
Results
IEE can accurately and robustly quantify causality
@@ Results
IEE can accurately capture the numerical behavior of dynamical causality in the two-node Logistic system, especially IntDC rather than ObsDC. In Fig. 2(A), we set $\beta_{xy}\equiv 0$ and let $\beta_{yx}$ increase from 00 to $0.3$. We observed a monotonic increase in $\text{IEE}[y\rightarrow x]$ (the red line with squares), while $\text{IEE}[x\rightarrow y]$ remained around zero (the blue line with dots). In Fig. 2(B), we set $\beta_{xy}\equiv 0.1$ and let $\beta_{yx}$ increase from 00 to $0.3$. $\text{IEE}[y\rightarrow x]$ exhibited a monotonic increase (the red line with squares), while $\text{IEE}[x\rightarrow y]$ stayed above zero (the blue line with dots). The gray dashed lines in Figs. 2(A-B) represent $0.01$ for reference. A detailed comparison with GC, TE and CCM can be found in figs. S1 and S2, which indicate that GC exhibited non-monotonic behavior, $\text{TE}[y\rightarrow x]$ slightly decreased with a false-negative issue, and $\text{CCM}[x\rightarrow y]$ had a false-positive problem. The mean values and standard deviations of the indices under 100 simulations are provided in tables S1-S4.
@@
Within suitable parameter ranges, IEE demonstrates stability and robustness. In Figs. 2(D-G), we validated the robustness of the IEE algorithm under various conditions, including different delayed lags $L$, the numbers of nearest neighbors $K$, lengths of time series $N$, and noise amplitudes $\sigma$. The causal strength should exhibit a descending trend with decreasing $\beta_{yx}$ from $0.15$ to $0.05$ through $0.125,0.1,0.075$. The experiments support that IEE maintains the correct order under different parameter values with relatively stable variations.
IEE can eliminate the effects of confounding variables
@@
IEE remains applicable and accurate even in presence of confounding variables. We simulated a three-node Logistic dynamics $(x_{t},y_{t},z_{t})$, where $z$ acted as a confounder (see details in SM). There were consistently non-zero causal effects from $z$ to $x$ and from $z$ to $y$, while $\beta_{xy}$ adjusted the causal strength from $x$ to $y$ (Fig. 2(C)). When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ was almost zero and significantly different from $\text{IEE}[z\rightarrow x]$ and $\text{IEE}[z\rightarrow y]$ (Fig. 2(C)); but conventional ObsDC indices such as GC/TE/CCM produced false-positive causality from $x$ to $y$ due to the presence of confounder $z$ (fig. S3). That is because an intervention on $x$, i.e. $\delta x$, will not induce $\delta y$ through $z$ under the IntDC framework. IEE, specially designed for measuring IntDC, remained unaffected by the confounder $z$ and yielded accurate results.
@@
Figure 2: Performance of IEE on the Logistic systems. (A) and (B) used the two-node Logistic system. The $\beta_{xy}$ and $\beta_{yx}$ control the coupling coefficients from $x$ to $y$ and from $y$ to $x$, respectively. (A) shows the mean values of IEE when $\beta_{xy}=0$ and $\beta_{yx}$ ranges from 00 to $0.3$ over 100 simulations. $\text{IEE}[y\rightarrow x]$ (the red line with squares) increases monotonically, while $\text{IEE}[x\rightarrow y]$ (the blue line with dots) stays around zero. (B) is under $\beta_{xy}=0.1$, in which $\text{IEE}[x\rightarrow y]$ (the blue line with dots) is significantly positive. The gray dashed lines in (A) and (B) represent the constant $0.01$ as reference. (C) shows the performance of IEE in the three-node Logistic system, where $z$ acts as a confounding variable. When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ is almost zero correctly. IEE designed for measuring IntDC is not affected by confounders, which usually leads to false positives. The results for $\beta_{xy}=0.5$ is shown for a comparison. The gray dashed line is $0.05$ for reference. (D-G) demonstrate the robustness of IEE under different delayed lags $L$, the numbers of nearest neighbors $K$, data lengths $N$, and noise amplitudes $\sigma$ in the two-node logistic system, respectively. The parameter $\beta_{xy}\equiv 0$, and $\beta_{yx}$ takes values of $0.15$ (red), $0.125$ (blue), $0.1$ (black), $0.075$ (green), and $0.05$ (magenta). Lines denote mean values under 50 simulations, and shaded areas represent the standard deviation. IEE maintains correct orders of IntDC under different situations with relatively stable variations.
IEE can reconstruct causal networks with quantified causal strength
@@
To verify the effectiveness of IEE in reconstructing causal networks, we calculated the Area Under Curve (AUC) values based on multiple simulated time series. The AUC value of IEE was $0.871\pm 0.008$, significantly higher than conventional ObsDC indices, such as GC with $0.707\pm 0.069$, TE with $0.816\pm 0.018$, and CCM with $0.837\pm 0.012$. The boxplots for AUC values are displayed in Fig. 3(B). AUCs for IEE exhibited a higher mean and a smaller variation compared to the others. Additionally, fig. S4 presents receiver operating characteristic (ROC) curves.
To show the capability for ranking the importance of causal influences, we conducted two tests. We first calculated the IEE from Node 1 to the other nine nodes. Consistently with the actual scenario, IEE displayed sequential decrease in IntDC originating from Node 1 across Nodes 2 to 10 (Fig. 3(C)). Then, we measured the IEE received by Node 7 from the other nine nodes. As expected, Nodes 1-6 exhibited ascending causal strengths on Node 7, whereas Nodes 8-10 had little influence on Node 7 (Fig. 3(D)). IEE accurately discerned the IntDCs, with Node 6 showing the strongest value. A comparison between IEE for the IntDC and the results from GC, TE, and CCM for the ObsDC can be found in figs. S5 and S6. GC and TE failed to rank the influence from Node 1 accurately (fig. S5), while CCM suffered from the false-positive causal detection from Node 8 to Node 7 (fig. S6).
Figure 3: Performance of IEE on the 10-node coupled Henon-map dynamics. (A) illustrates the network structure of the dynamics. There are directed causality from each node to its subsequent node. (B) presents a boxplot of the Area Under Curve (AUC) values for the network reconstruction. IEE outperforms GC, TE, and CCM significantly in accuracy and stability. (C) demonstrates the IEE from Node 1 to the other nine nodes, while (D) shows the IEE received by Node 7 from the other nodes. IEE accurately captures the interventional flow and IntDC in cascade topology.
IEE can quantify IntDC effectively without requiring interventions or perturbations
@@
The chNN comprises an output variable $\bm{x}=(x_{1},x_{2},\dots,x_{N_{\text{node}}})^{T}$ and two internal variables, with $N_{\text{node}}$ representing the number of neurons in the network. We chose $N_{\text{node}}=10$ in simulations. Details on the dynamics can be found in SM. We denoted the observed time-series data from the stationary system without intervention as $\bm{x}_{\text{obs}}$. When the neuron $i$ was removed (set as a constant zero), the perturbed data were denoted as $\bm{x}_{\text{per}}^{(i)}$. Removing one node was taken as an intervention to the system. The stationary probability density functions of the $j$th neuron before and after the intervention were represented by $p_{j}(\bm{x}_{\text{obs}})$ and $p_{j}(\bm{x}_{\text{per}}^{(i)})$, respectively. We used the Kullback-Leibler divergence (KLD), i.e.
@@
to quantify the true influence of the intervention, i.e. IntDC, from $i$ to $j$. By conducting $100$ randomly simulated chNNs with interventions (10 removed nodes for each chNN), we recorded $1000$ KLD values and compared them with the results of IEE (calculated from $\bm{x}_{\text{obs}}$ solely). In Fig. 4(A), gray dots are the $1000$ samples of causal edges between neurons, while the red line stands for the linear regression. We observed that $\text{IEE}[x_{i}\rightarrow x_{j}]$ exhibited a positive linear correlation with the KLD $D_{ij}$ (with $R^{2}=0.869$). Results for GC (with $R^{2}=0.572$), TE (with $R^{2}=0.912$), and CCM (with $R^{2}=0.622$) are shown in Figs. 4(B-D), which imply that GC and CCM cannot linearly reflect KLD. Furthermore, the violin plots of cosine similarity $S_{c}$ (see SM) between KLD and the four causal indices on $100$ chNNs are displayed in Fig. 4(E). IEE ($0.904\pm 0.024$) had the highest mean similarity with KLD, compared to GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$). These results indicate that IEE can accurately quantify IntDC from $\bm{x}_{\text{obs}}$ alone, alleviating the need for additional interventions to the dynamics.
@@
Figure 4: The performance of IEE when measuring the influence of perturbations on the chaotic neural networks (chNNs). The Kullback-Leibler divergence (KLD) is utilized to measure the true influence or IntDC between neurons. In (A-D) , $1000$ samples for different perturbations on chNNs are represented by gray dots, along with the linear regressions (the red lines) between KLD and IEE/GC/TE/CCM, respectively. (E) displays the violin plots of cosine similarity between KLD and IEE/GC/TE/CCM. The mean values and standard variations of the cosine similarity are IEE ($0.904\pm 0.024$), GC ($0.711\pm 0.076$), TE ($0.888\pm 0.022$), and CCM ($0.687\pm 0.033$). IEE can linearly reflect the KLD with high similarity, indicating its effectiveness in accurately quantifying IntDC without additional perturbed data on chNNs.
Application of IEE in inferring neural connectomes of C. elegans
@@
We applied the IEE criterion to infer the neural connectomes of Caenorhabditis elegans (C. elegans), a model organism known for its comprehensively studied nervous system  (?). Calcium fluorescence imaging time series data from $31$ neurons with specific functions of a freely moving C. elegans were collected from Kato et al.  (?) (see Figs. 5(A-B) and fig. S7) . The Fig. 5(C) shows the cosine similarity matrix between different neurons. We clustered the neurons into $7$ clusters after data preprocessing (see details in Supplementary Text and table S5). Figure 5(D) presents the ground truth of neural connectomes in a directed graph, which was detected by electron microscopy in references  (?, ?, ?). We calculated the IEE between different clusters to infer the IntDC network. The ROC curve of IEE is shown in Fig. 5(F) with an AUC value $0.882$. The maximum Youden index, maximum concordance probability, and minimum distance to the point $(0,1)$ gave the same optimal operating point (OOP). At the OOP, Fig. 5(E) displays the inferred connectomic network by IEE. Comparisons with other ObsDC indices, i.e. GC/TE/CCM, revealed that IEE exhibited the highest AUC value, best OOP, largest similarity to the ground truth, lowest false positive at the OOP, and lowest false negative at the OOP (see Table 2, figs. S8-S10, and table S6). These findings demonstrate the efficacy of IEE in reconstructing the neural connectomes of C. elegans.
Figure 5: Application of the IEE on inferring the neural connectomes of C. elegans. (A) is an illustration of C. elegans. (B) displays the calcium imaging time series of $31$ key neurons, with their names listed on the left and colors representing different clusters. (C) is the cosine similarity matrix between neurons, aiding the clustering process. (D) is the true connectomes between $7$ clusters, while (E) is the inferred causal network at the the optimal operating point (OOP) determined by IEE. The red edge represents a false positive, and the black dashed edges represent false negatives. (F) shows the ROC curve (the blue line) for IEE. The same OOP is obtained under three criteria, i.e. the maximum Youden index (the blue circle), the maximum concordance probability (the green dot), and the minimum distance to the point $(0,1)$ (the red cross). The shaded area around the ROC curve represents the $95\%$ confidence interval obtained through bootstrapping. The AUC value of IEE is $0.882$.
Application of IEE to COVID-19 transmission in Japan

### v2

@@ 2.3 Interventional embedding entropy
We introduce the IEE criterion, specifically designed for quantifying IntDC solely from the observational data in the delay-embedding space.
Definition 3 (Interventional embedding entropy, IEE).
The IEE criterion is proposed to measure the average information retention or reconstructability from the effect $\delta\bm{Y}_{t+1}$ to the causation $\delta\bm{X}_{t}$, formulated as
$\mathrm{IEE}[x\rightarrow y]:=\mathrm{CMI}(\delta\bm{X}_{t},\delta\bm{Y}_{t+1}|\bm{Y}_{t+1}).$ (7)
In Eq. (7), CMI is the conditional mutual information  (?), i.e.
@@
IEE provides is a quantitative measure of causal strength that is comparable across different edges within the same dynamical system. When dealing with data, the $\delta\bm{Y}_{t+1}$ is estimated by the distance from $\bm{Y}_{t+1}$ to its neighbored points $\bm{Y}_{t_{k}+1}$ in the embedding space, and $\delta\bm{X}_{t}$ is calculated by $\bm{X}_{t_{k}}$ with the same time label $t_{k}$ (Figs. 1(E) and (F)). Thus, Eq. (7) can be numerically approximated only from the observational data, especially for non-intervention systems. Table 1 presents the numerical algorithm for computing IEE, whose detailed descriptions and computational procedures are provided in the Supplementary Text.
2.4 Connections with statistical causal models and assumptions of ConDC/IntDC
The concept of dynamical causality in Eq. (1) develops from traditional structural causal models, which typically studies structure equations of the form $Y=f(X,\varepsilon,\theta)$, where $X$ and $Y$ are random variable, $f$ is a function parameterized by $\theta$, and $\varepsilon$ represents noise or residue terms. In  (?), the framework of ConDC is systematically studied with three key generalizations beyond classical structural equation modeling: (i) Temporal information is fully considered by dynamical systems, ensuring that causes precede effects in time; (ii) Feedback or causal loops between variables are permitted under the consideration of time delays, relaxing the restriction of directed acyclic graphs (DAGs); (iii) Quantitative measures of causal strength are provided beyond qualitative identification, enabling direct comparisons of causality between different variables. The IntDC proposed in the study further integrates the concept of interventions and potential outcomes within the dynamical framework (?). IntDC aims to quantify causal strengths at the interventional level directly from observational data (without the need for additional interventions to the system, which are often impractical or infeasible in real-world datasets).
The following assumptions (A1)-(A8) clarify key connections between dynamical causality and statistical causal inference:
(A1) Temporal Order assumption: Causal information is embedded in time-series data and causes must precede their effects;
(A2) Causal Sufficiency (Unconfoundedness) assumption: All common drivers (confounders) are observed such that the causal graph accurately represents relationships among the observed variables;
(A3) Causal Markov assumption: Each variable $x_{i}$ is independent of its nondescendants given its direct causes (parents);
(A4) Faithfulness assumption: The causal graph structure can represent the conditional independence contained in the joint probability density;
(A5) Stationarity assumption: The dynamical system evolves into a stable attractive manifold, allowing stable causal relations to be measured from observational data;
(A6) Intervention Ignorability assumption: Given the system dynamics Eq. (1), initial conditions $\bm{x}_{0}$ and noise $\bm{\varepsilon}_{t}$, the intervention on variable $x^{(j)}$ at time $t-k$ is independent of the potential outcome of $x^{(i)}$ at time $t+1$, i.e. $x^{(i)}_{t+1}\text{$\perp\!\!\!\perp$}\delta x^{(j)}_{t-k}|(\bm{f},\bm{x}_{0},\bm{\varepsilon}_{t})$ for $k=1,2,\dots,p$;
(A7) Stable Infinitesimal Intervention assumption: The average information retention from $x^{(j)}$ to $x^{(i)}$ in dynamics Eq. (1) remains stable under sufficiently small interventions $\delta x^{(j)}$;
(A8) Consistency assumption: The outcome $\tilde{x}^{(i)}_{t+1}$ in Eq. (1) can be precisely determined under known dynamics $(\bm{f},\bm{x}_{0},\bm{\varepsilon}_{t})$ and a fixed intervention $\delta x^{(j)}_{t-k}$.
In practice, (A2) can be relaxed if the primary goal is to quantify the total causal strength (including both direct and indirect causality) between observed variables. In the Results section, we demonstrate that the IEE criterion is capable of mitigating the impact of confounding variables, thereby providing robust causal estimates even when some confounders are unobserved. Some of these assumptions for causal discovery from observational time-series data have also been discussed by Runge et al. ( (?, ?)).
GC, TE, and CCM are all criteria for ConDC, which estimate causal strengths by fitting constructive dynamical models (i.e. Eq. (2) in the time-series space or Eq. (5) in the delay-embedding space). GC employs linear vector regression to fit $\bm{f}$ in Eq. (2). TE extends GC to nonlinear cases by quantifying information transfer between variables. CCM constructs a local-linear cross-mapping $\bm{F}$ in Eq. (5) by the time-delayed embedding to capture causality. For a detailed comparison of these ConDC methods, readers are referred to Table 1 in  (?).
3 Results
3.1 IEE captures the numerical behavior of causality in Logistic systems robustly
@@
IEE can capture the numerical behavior of IntDC in the two-node Logistic system. When we set $\beta_{xy}\equiv 0$ and let $\beta_{yx}$ increase from 0 to 0.30.30.3. we observed a monotonic increase in $\text{IEE}[y\rightarrow x]$ (the red line with squares in Fig. 2(A)), while $\text{IEE}[x\rightarrow y]$ remained around zero (the blue line with dots in Fig. 2(A)). When we set $\beta_{xy}\equiv 0.1$ and let $\beta_{yx}$ increase from 0 to 0.30.30.3, $\text{IEE}[y\rightarrow x]$ exhibited a monotonic increase (the red line with squares in Fig. 2(B)), while $\text{IEE}[x\rightarrow y]$ stayed above zero (the blue line with dots in Fig. 2(B)). The gray dashed lines in Figs. 2(A-B) represent 0.010.010.01 for reference. A detailed comparison with GC, TE and CCM is presented in figs.S1 and S2. The mean values and standard deviations of the indices over 100 simulation runs are provided in tables S1-S4. Fig. S1(b) and fig. S2(b) indicate that GC exhibited non-monotonic behavior. $\text{TE}[y\rightarrow x]$ demonstrated a decreasing trend when $\beta_{yx}$ increased, specifically for $\beta_{yx}\in(0,0.02]$ in fig. S1(c)/table S1 and $\beta_{yx}\in(0,0.03]$ in fig. S2(c)/table S3, which may lead to false-negative issues (ignorance of existence causality). $\text{CCM}[x\rightarrow y]$ remained significantly different from zero even when $\beta_{xy}=0$ (in the range $\beta_{yx}\in[0.15,0.3]$ in fig. S1(d)/table S2). $\text{CCM}[y\rightarrow x]$ was also significantly different from zero when $\beta_{yx}=0$ (with $\beta_{xy}=0.1$ in fig. S2(d)/table S3). These results suggest that CCM may suffer from the false-positive problem (incorrect identification of non-existence causality).
@@
Within suitable parameter ranges, IEE demonstrates stability and robustness. We validated the robustness of the IEE algorithm under various conditions, including different delayed lags $L$, the numbers of nearest neighbors $K$, lengths of time series $N$, and noise standard deviations $\sigma$. The causal strengths measured by IEE exhibited a consistent descending trend as $\beta_{yx}$ decreased from 0.150.150.15 to 0.050.050.05 through 0.125,0.1,0.0750.125,0.1,0.0750.125 , 0.1 , 0.075, as expected (Figs. 2(D-G)). These results support that IEE reliably preserves the correct ranking of causal strengths across different parameter settings with relatively stable variance.
3.2 IEE eliminates the influence of confounding variables
@@
IEE remains applicable and accurate even in presence of confounding variables. We simulated a three-node Logistic dynamics $(x_{t},y_{t},z_{t})$, where $z$ acted as a confounder (see details in SM). There were consistently non-zero causal effects from $z$ to $x$ and from $z$ to $y$, while $\beta_{xy}$ adjusted the causal strength from $x$ to $y$ (Fig. 2(C)). When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ was almost zero and significantly different from $\text{IEE}[z\rightarrow x]$ and $\text{IEE}[z\rightarrow y]$ (Fig. 2(C)); but conventional ConDC indices such as GC/TE/CCM produced false-positive causality from $x$ to $y$ due to the presence of confounder $z$ (fig. S3). According to the IntDC framework, an intervention on $x$ (i.e. $\delta x$) will not induce an indirect change in $y$ (i.e. $\delta y$) through $z$; in other words, there is no intervention-induced pathway through $x\leftarrow z\rightarrow y$. However, traditional ConDC indices by fitting dynamical models may falsely infer an association between $x$ and $y$ due to the influence of the confounder $z$. IEE, specially designed for measuring IntDC, remained unaffected by the confounder $z$ and yielded accurate results.
@@
Figure 2: Performance of IEE on Logistic systems. (A) and (B) used the two-node Logistic system. The $\beta_{xy}$ and $\beta_{yx}$ control the coupling coefficients from $x$ to $y$ and from $y$ to $x$, respectively. (A) shows the mean values of IEE when $\beta_{xy}=0$ and $\beta_{yx}$ ranges from 0 to 0.30.30.3 over 100 simulations. $\text{IEE}[y\rightarrow x]$ (the red line with squares) increases monotonically, while $\text{IEE}[x\rightarrow y]$ (the blue line with dots) stays around zero. (B) is under $\beta_{xy}=0.1$, in which $\text{IEE}[x\rightarrow y]$ (the blue line with dots) is significantly positive. The gray dashed lines in (A) and (B) represent the constant 0.010.010.01 as reference. (C) shows the performance of IEE in the three-node Logistic system, where $z$ acts as a confounding variable. When $\beta_{xy}=0$, $\text{IEE}[x\rightarrow y]$ is almost zero correctly. IEE designed for measuring IntDC is not affected by confounders, which usually leads to false positives in ConDC. The results for $\beta_{xy}=0.5$ is shown for a comparison. The gray dashed line is 0.050.050.05 for reference. (D-G) demonstrate the robustness of IEE under different delayed lags $L$, the numbers of nearest neighbors $K$, data lengths $N$, and noise standard deviations $\sigma$ in the two-node logistic system, respectively. The parameter $\beta_{xy}\equiv 0$, and $\beta_{yx}$ takes values of 0.150.150.15 (red), 0.1250.1250.125 (blue), 0.10.10.1 (black), 0.0750.0750.075 (green), and 0.050.050.05 (magenta). Lines denote mean values under 50 simulations, and shaded areas represent the standard deviation. IEE consistently preserves the correct ranking of IntDC across various settings, with relatively stable variance.
3.3 IEE enables causal network reconstruction with quantification of causal strengths
@@
To verify the effectiveness of IEE in reconstructing causal networks, we calculated the Area Under Curve (AUC) values based on multiple simulated time series. The AUC value (mean $\pm$ standard deviation) of IEE was 0.871±0.0080.871$\pm$ 0.0080.871 ± 0.008, significantly higher than conventional ConDC indices, such as GC with 0.707±0.0690.707$\pm$ 0.0690.707 ± 0.069, TE with 0.816±0.0180.816$\pm$ 0.0180.816 ± 0.018, and CCM with 0.837±0.0120.837$\pm$ 0.0120.837 ± 0.012 (boxplots in Fig. 3(B)). The AUCs of IEE exhibited both a higher mean and a lower variance relative to other approaches. Additionally, receiver operating characteristic (ROC) curves for the four indices were presented in fig. S4.
To show the capability for ranking the importance of causal influences, we conducted two tests. We first calculated the IEE from Node 1 to the other nine nodes. Consistently with the actual scenario, IEE displayed sequential decrease in IntDC originating from Node 1 across Nodes 2 to 10 (Fig. 3(C)). Then, we measured the IEE received by Node 7 from the other nine nodes. As expected, Nodes 1-6 exhibited ascending causal strengths on Node 7, whereas Nodes 8-10 had little influence on Node 7 (Fig. 3(D)). IEE accurately discerned the IntDCs, with Node 6 showing the strongest value. A comparison showed that GC and TE failed to rank the influence from Node 1 accurately in the first test (fig. S5), while CCM suffered from the false-positive causal detection from Node 8 to Node 7 in the second test (fig. S6).
Figure 3: Performance of IEE on the 10-node coupled Henon-map dynamics. (A) illustrates the network structure of the dynamics. There are directed causality from each node to its subsequent node. (B) presents a boxplot of the Area Under Curve (AUC) values for the network reconstruction. IEE (0.871±0.0080.871$\pm$ 0.0080.871 ± 0.008) outperforms GC (0.707±0.0690.707$\pm$ 0.0690.707 ± 0.069), TE (0.816±0.0180.816$\pm$ 0.0180.816 ± 0.018), and CCM (0.837±0.0120.837$\pm$ 0.0120.837 ± 0.012) significantly in accuracy and stability. (C) demonstrates the IEE from Node 1 to the other nine nodes, while (D) shows the IEE received by Node 7 from the other nodes. IEE accurately captures the interventional information flow and IntDC in the cascade topological network.
3.4 IEE quantifies IntDC without requiring additional experimental interventions
@@
The chNN comprises an output variable $\linebreak\bm{x}=(x_{1},x_{2},\dots,x_{N_{\text{node}}})^{T}$ and two internal variables, with $N_{\text{node}}$ representing the number of neurons in the network. We chose $N_{\text{node}}=10$ in simulations. Details on the dynamics can be found in SM. We denoted the observed time-series data from the stationary system without intervention as $\bm{x}_{\text{obs}}$. When the neuron $i$ was removed (set as a constant zero), the perturbed data were denoted as $\bm{x}_{\text{per}}^{(i)}$. Removing one node was taken as an intervention to the system. The stationary probability density functions of the $j$th neuron before and after the intervention were represented by $p_{j}(\bm{x}_{\text{obs}})$ and $p_{j}(\bm{x}_{\text{per}}^{(i)})$, respectively. We used the Kullback-Leibler divergence (KLD), i.e.
@@
to quantify the true influence of the intervention, i.e. ground truth of the IntDC, from $i$ to $j$. Liang’s Information flow adopts a similar intervention concept by treating a variable as fixed  (?, ?). By conducting 100100100 randomly simulated chNNs with interventions (in each simulation, one node is removed from each chNN with $N_{\text{node}}$ nodes), we recorded 100010001000 KLD values and compared them with the results of IEE (calculated from $\bm{x}_{\text{obs}}$ solely). We observed that $\text{IEE}[x_{i}\rightarrow x_{j}]$ exhibited a positive linear correlation with the KLD $D_{ij}$ (with $R^{2}=0.869$) in Fig. 4(A), where gray dots are the 100010001000 samples of causal edges between neurons and the red line stands for the linear regression. Results for GC (with $R^{2}=0.572$), TE (with $R^{2}=0.912$), and CCM (with $R^{2}=0.622$) indicated that GC and CCM could not linearly reflect KLD (Figs. 4(B-D)). Furthermore, we drew the violin plots of cosine similarity $S_{c}$ (see SM) between KLD and the four causal indices on 100100100 chNNs (Fig. 4(E)). IEE (0.904±0.0240.904$\pm$ 0.0240.904 ± 0.024) had the highest mean similarity with KLD, compared to GC (0.711±0.0760.711$\pm$ 0.0760.711 ± 0.076), TE (0.888±0.0220.888$\pm$ 0.0220.888 ± 0.022), and CCM (0.687±0.0330.687$\pm$ 0.0330.687 ± 0.033). These results present that IEE can accurately quantify IntDC from $\bm{x}_{\text{obs}}$ alone, alleviating the need for additional interventions to the dynamics.
@@
Figure 4: The performance of IEE when measuring the influence of perturbations on the chaotic neural networks (chNNs). The Kullback-Leibler divergence (KLD) is utilized to measure the true influence or IntDC between neurons. In (A-D) , 100010001000 samples for different perturbations on chNNs are represented by gray dots, along with the linear regressions (the red lines) between KLD and IEE/GC/TE/CCM, respectively. (E) displays the violin plots of cosine similarity between KLD and IEE/GC/TE/CCM. The mean values and standard deviations of the cosine similarity are IEE (0.904±0.0240.904$\pm$ 0.0240.904 ± 0.024), GC (0.711±0.0760.711$\pm$ 0.0760.711 ± 0.076), TE (0.888±0.0220.888$\pm$ 0.0220.888 ± 0.022), and CCM (0.687±0.0330.687$\pm$ 0.0330.687 ± 0.033). IEE can linearly reflect the KLD with high similarity, indicating its effectiveness in accurately quantifying IntDC without additional perturbed data on chNNs.
3.5 Application of IEE in inferring neural connectomes of C. elegans
@@
We applied the IEE criterion to infer the neural connectomes of Caenorhabditis elegans (C. elegans), a model organism known for its comprehensively studied nervous system  (?). Calcium fluorescence imaging time series data from 313131 neurons with specific functions of a freely moving C. elegans were collected from Kato et al.  (?) (see Figs. 5(A-B) and fig. S7). The cosine similarity between different neurons were calculated and the neurons were clustered into 777 clusters after data preprocessing (see Fig. 5(C), Supplementary Text and table S5). We presented the ground truth of neural connectomes between clusters in a directed graph (Fig. 5(D)), which was detected by electron microscopy  (?, ?, ?). The IEE values between different clusters were calculated to infer the IntDC network. The ROC curve of IEE got a high AUC value 0.8820.8820.882 (Fig. 5(F)). By maximum Youden index, maximum concordance probability, and minimum distance to the point (0,1)(0,1)( 0 , 1 ), IEE gave the same optimal operating point (OOP). At the OOP, IEE provided an inferred connectomic network with only 1 false positive and 4 false negatives (Fig. 5(E)). We also compared IEE with other ConDC indices, i.e. GC/TE/CCM, which revealed that IEE exhibited the highest AUC value, best OOP, largest similarity to the ground truth, lowest false positive at the OOP, and lowest false negative at the OOP (see Table 2, figs. S8-S10, and table S6).
Additionally, we inferred the connectomic network using the PCMCI algorithm with two approaches: the partial correlation (PCMCI-ParCorr) and CMI test (PCMCI-CMI). PCMCI used an independence test with the only parameter being the significance level $\alpha_{\text{PC}}$. The false positive edges for PCMCI-ParCorr ($\alpha_{\text{PC}}=0.05$) and PCMCI-CMI ($\alpha_{\text{PC}}=0.01$) are 555 and 666, respectively. Both approaches have 444 false negative edges. A detailed description of the procedure and results can be found in the Supplementary Text, figs. S10(e-f) and table S6.
These findings demonstrate the efficacy of IEE in reconstructing the neural connectomes of C. elegans.
[…]
3.6 Application of IEE to COVID-19 transmission in Japan